total of 698 tickets were sold for an event. Three times as many gold (g) tickets were sold as silver (s) tickets. The total number of gold and silver tickets sold was twice the number of bronze (b) tickets sold. Set up equations, don't actually solve.
Answer by mananth(16946)

let number of silver tickets be s
gold = g = 3s
..
bronze = (3s+s)= 2b
/2
b=4s/2 = 2s
...
Total number of tickets = 698
...
s+g+b
= s+3s+2s =698
=6s=698
...
m.ananth@hotmail.ca